Solve the following system of equations. (Enter your answers as a comma-separated list. If there are infinitely many solutions, enter a parametric solution using t and/or s. If there is no solution, enter NONE.) x1 + 2x2 + 6x3 2 x1 + x2 + 3x3 = 1
Answer #1

multiply equation 2 with 2=2x1+2x2+6x3=2-------(3)

x1+2x2+6x3=2-------(1)

subtract 3 from 1= x1=0

substitute value of x1 in equation 1,2

2x2+6x3=2

x2+3x3=1

both equations are same so there are infinte number of solution

let x2=t

x2+3x3=1

t+3x3=1

x3=(1-t)/3

(x1,x2,x3)=(0,t,(1-t)/3)